0.5N titrant is used to reach the equivalence point with 800mL of an acetic acid solution with a pH of 3. What volume of titrant is needed to reach equivalence with the acetic acid solution?
1
Expert's answer
2020-04-02T08:55:50-0400
  • For the beginning: "pH = -lg[H^+] \to" [H+] = 10-3
  • Because acetic acid is a weak acid, therefore this formula should be applied: "[H^+] = \sqrt (Ka•C(acid))" , then, the concentration will be calculated: "C(acid) = [H^+]^2 \/ Ka" = (10-3)2 / 1.8•10-5 = 0.06 M
  • Since the conditions don't indicate what kind of titrant was applied, normality will be equal to molarity. At the equivalence point, equality is observed for reacting substances: "C(acid) \u2022 V(acid) = C(titrant) \u2022 V(titrant)". Then: "V(titrant) = (C(acid)\u2022V(acid)) \/ C(titrant)" = (0.06•0.8) / 0.5 = 0.096 L = 96 ml
  • Answer: V(titrant)=96 ml.